OpenStudy (anonymous):

need help with these 2 problems: 1. a horizontal force acting on a 3.0 kg object produces a horizontal acceleration of 6.0 m/s^2.find the magnitude of the net force acting on the object? 2. a force gives acceleration of 3m/s^2 to a 1kg body.if the same force accelerates another body at 9m/s^2, what is the mass of the second body?

OpenStudy (anonymous):

1. 18 N 2. 1/3 kg...

OpenStudy (anonymous):

can u show me the formula on how did u get those answers?

OpenStudy (anonymous):

\[\Sigma F=ma\] for a: \[\Sigma F=3\times6=12 N\]

OpenStudy (anonymous):

how about the number 2 question?

OpenStudy (anonymous):

Find sigma F first with the same equation. What's sigma F?

OpenStudy (anonymous):

so u mean the one u gave formula to get \[\Sigma F\]?

OpenStudy (anonymous):

Yes, find sigma F now

OpenStudy (anonymous):

1/9 or .111 is the answer?

OpenStudy (anonymous):

sigma F=3*1=3 N m=F/a=3/9=1/3

OpenStudy (anonymous):

Thanks..